OpenStudy (anonymous):

Can anyone help me solve this? Simplify (–3)2 + (–4 + 2)3 ÷ 4 – 3. A. –14 B. –8 C. 4 D. 11

OpenStudy (anonymous):

first you know the form PEMDAS right??

OpenStudy (anonymous):

first you have to do the perenthesis first so (-3)2 + (-4 + 2)3 / 4 - 3 = (-6 + -6) / 4 - 3 = (-12) / 4 - 3 = -3 -3 = -6

OpenStudy (anonymous):

are you sure that the questions is right??

OpenStudy (anonymous):

That's what I got! :( But it doesn't seem right.. I don't know!! :(

OpenStudy (anonymous):

D

OpenStudy (anonymous):

how you get D??

OpenStudy (anonymous):

I did the work

OpenStudy (anonymous):

show it on the chat

OpenStudy (anonymous):

Can you please explain @Oscar123Flores

OpenStudy (anonymous):

ok i might have got it. wait

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

check your questions again if it is right

OpenStudy (anonymous):

Simplify (–3)2 + (–4 + 2)3 ÷ 4 – 3. A. –14 B. –8 C. 4 D. 11

OpenStudy (anonymous):

they are like this...

OpenStudy (anonymous):

Express 3/4 (2-4) as a single fraction. 3/4 (2-4) = (3 (2-4))/4: -3×2+(3 (2-4))/4-3 Subtract 4 from 2. 2-4 = -2: -3×2+(3×-2)/4-3 In (-2×3)/4, the numbers -2 in the numerator and 4 in the denominator have gcd greater than one. The gcd of -2 and 4 is 2, so (-2×3)/4 = ((2 (-1)) 3)/(2×2) = 2/2×(-3)/2 = (-3)/2: -3×2+(3×-1)/2-3 Multiply -3 and 2 together. -3×2 = -6: -6-3/2-3 Put the fractions in -6-3/2-3 over a common denominator. Put -6-3/2-3 over the common denominator 2. -6-3/2-3 = (2 (-6))/2-3/2+(2 (-3))/2: (-6×2)/2-3/2-(3×2)/2 Multiply -6 and 2 together. -6×2 = -12: -12/2-3/2-(3×2)/2 Multiply -3 and 2 together. -3×2 = -6: (-12)/2-3/2+-6/2 Add the fractions over a common denominator to a single fraction. (-12)/2-(3)/2-(6)/2 = (-12-3-6)/2: (-12-3-6)/2 Group the negative terms in -12-3-6 together and factor out the minus sign. -12-3-6 = -(12+3+6): -(12+3+6)/2 Evaluate 12+3+6 using long addition. | 1 | | 1 | 2 | | 6 + | | 3 | 2 | 1: Answer: | | (-21)/2

OpenStudy (anonymous):

that is 11.5 not 11 exatly it did not say to round it said to simplify.

OpenStudy (anonymous):

trust me

OpenStudy (anonymous):

I know this things

OpenStudy (anonymous):

and it does not look like you followed the rule of PEMDAS. P = Perenthesis E = Expression M = Multilication D = Division A = Addition S = Subtraction

OpenStudy (anonymous):

or E = Exponents

OpenStudy (anonymous):

I know what pemdas is

OpenStudy (anonymous):

wait is it suppose to be (-3) ^ 2??

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

wow thats why you made us think that it was multiplication

OpenStudy (anonymous):

write the whole equation like the way i wrote it

OpenStudy (anonymous):

I am sorry!

OpenStudy (anonymous):

with a ^ for exponents

OpenStudy (anonymous):

(-3)^2 + (-4+2)^3 / 4-3

OpenStudy (anonymous):

i think it is b because (–3)^2 + (–4 + 2)^3 ÷ 4 – 3. 9 + -8 / 4 - 3 9 + -2 - 3 9 - 5 = 4

OpenStudy (anonymous):

are you good??

OpenStudy (anonymous):

Yes!! THank YOU!!